OpenStudy (anonymous):

For the real-valued functions f(x)=sqrt(5x+20) and g(x)=x-2 , find the composition f o g(x) and specify its domain using interval notation.

OpenStudy (anonymous):

f(g(x) = sqrt(5(x-2)+20) = sqrt(5x-10+20) = sqrt(5x+10) 5x+10>=0 since sqrt cannot contain negative numbers so 5x>-10 x>-2 so domain is [-2,infinite) and range [0,infinte)

OpenStudy (anonymous):

f(x)=sqrt(5x-10)

OpenStudy (anonymous):

Would domain be [2,infinity) ?

OpenStudy (anonymous):

It is from [-2,infinity), can you please be clear with your question

OpenStudy (anonymous):

Find the domain of the function f(x)=sqrt(5x-10) I typed f(x)=sqrt(5x+10) in the first question. Sorry for the confusion.